OpenStudy (anonymous):

Could I get some help with this question please:) Demand function 128 ------ +2 x^2 Quality demand x = 4 Find the price elasticity of demand for the demand function at the indicated x-value.

OpenStudy (anonymous):

2^7 = 128; (1/2)^x=128; (1/2)^-7 = 2^7; x = - 7 (4/9)^x = 81/16 = (9/4)^2; x = - 2 (1/8)^x = 1/16 = (1/8)/2 = (1/2^3)/2 = 1/2^4); x = 4/3 ( the cube root of 8^4)